OpenStudy (anonymous):

Find the solutions of the inequality. | x - 5 | ≥ 3

OpenStudy (anonymous):

You need to have two solution right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x-5=3 and x-5=-3 +5 +5 +5 +5 x=8 x=2 (8,2) are the solution